For what value of k,2x+ky=6 has equal values of x and y for its solution?

Grade:12

1 Answers

Arun
5 years ago
Given,
x=y
Now,
2x + ky = 6
or,ky = 6 - 2x
or,ky = 6 - 2y [as x=y]
or, k = 6 - 2y / y
So,
As the value of k depends on value of y.There will be different values of k for the different values of y